# MetaGS: A Meta-Learned Gaussian-Phong Model for Out-of-Distribution 3D Scene Relighting

### v2

@@ 1 Introduction
3D scene relighting generates novel lighting effects that interact with the observed 3D environment, with recent advances in learning-based volume rendering offering effective solutions [20, 35, 22, 39, 23, 21, 3, 29]. A typical approach captures a substantial number of multi-view images of a scene under individual lighting conditions, then trains a NeRF or 3D Gaussian splatting (3DGS) model to generalize to new point light positions [27, 12, 8, 40, 7]. Notably, such idealized data requirements are often impractical in real-world scenarios. To simulate a more realistic capturing process, we build on prior work using a One Light At a Time (OLAT) setup [33, 13, 1], where training data is collected with a moving point light and a moving camera. Recent studies have further explored low-cost OLAT capture using a smartphone flashlight as the moving light source [5].
@@
Figure 1: Preliminary results: NRHints [33] and GS3 [1] struggle with out-of-distribution light positions (as specified in Section. 2), as the blurry novel view synthesis results and erroneous normal predictions indicate. GS3 [1] does not require well-defined surface normals and does not directly output the normal image.
@@
Furthermore, MetaGS integrates simple yet fundamental physical priors into Gaussian splatting by incorporating a learnable Blinn-Phong reflection model [2]. This approach effectively decouples different shading components—diffuse, specular, and ambient—leading to a more physically grounded understanding of the interactions between objects and lighting in the training data. This design draws inspiration from prior research [38, 27, 12] indicating that modular and compositional representations enhance generalization capabilities. The decoupled rendering improves the model’s ability to generalize to novel lighting conditions, allowing it to independently adjust each shading component according to the specific lighting scenario.
@@ 2 Preliminaries
We summarize existing OLAT relighting methods in Table LABEL:tab:olat_methods. In the preliminary experiments, we evaluate NRHints [33], the prior art in NeRF-based relighting models. We observe that it underperforms in testing scenarios with out-of-distribution (OOD) lighting positions, a feature that we believe should be crucial for relighting methods. Specifically, in Figure 1, the lighting in the training set is arranged on one side of the hemisphere, while the lighting in the test set is arranged on the opposite side (cameras are located on both sides). In such cases, NRHints fails to generate reasonable rendering results, likely due to its implicit modeling of shadows and specular reflections. We additionally evaluate a concurrent 3DGS-based approach [1] and observe similar degraded results under the OOD relighting setup. These findings highlight that understanding intrinsic illumination properties under arbitrary lighting variations remains a challenging task, as the model tends to overfit to perspective-constrained observations and fail to leverage enough physical principles when dealing with unseen lighting distribution.
@@
[34] [5] [6] [33] [1] [11] Ours
@@ 3.1 Differentiable Phong Model
Our MetaGS method leverages simple yet generalizable physical priors from the Blinn-Phong model [18], which captures three fundamental components of light transport: ambient, diffuse, and specular reflections. The core idea is to disentangle these illumination components by learning the interactions between (i) the normal vectors of the Gaussian points, (ii) the viewing directions, and (iii) the ray direction from the point light.
@@
where $I$ denotes the light emitted intensity, which is a global learnable parameter, $\mathbf{l}$ denotes the point-to-light normalized vector, and $\mathbf{h}=\frac{\mathbf{v}+\mathbf{l}}{||\mathbf{v}+\mathbf{l}||}$ denotes the bisector of the point-to-camera normalized vector $\mathbf{v}$ and $\mathbf{l}$. We model the coefficients $k_{d}$ and $k_{s}$ implicitly, where $k_{s}$ is multiplied by the RGB color of the point light.
@@
We calculate shadow effects using a BVH-based ray tracing method [7]. Similar to the camera-to-point accumulated transmittance, the received light intensity $T_{i}^{\text{light}}$ represents the total transmittance of light along the light-to-point ray. It is therefore affected by the opacity of the surfaces encountered along this path. For each point, we determine light visibility by tracing a ray from the Gaussian’s center to the light source. Since only the diffuse and specular colors are influenced by incident light intensity, we update the color of each point by incorporating the light visibility factor into the diffuse and specular terms from Eq. (1):
@@ 3.2 Meta-Learned Gaussian Relighting
To address this, we introduce a meta-learning framework based on bilevel optimization, which has been shown to effectively bridge the distribution shift between the training and testing domains, facilitating the generalization of optimized variables to unseen scenarios [4]. This training strategy can also promote coherent geometry learning in complex OLAT tasks, as shown in the ablation study. In MetaGS, the intuition of incorporating bilevel gradient update is to mitigate overfitting to specific light conditions by explicitly simulating test samples with OOD light sources during each gradient update, thereby improving the model’s ability to generalize to varied lighting scenarios.
@@
[…]
@@
9:    $\{\theta\}\leftarrow\{\theta\}-\beta\sum_{i=1}^{m}\nabla_{\theta}\mathcal{L}_{\text{final}}(\theta_{i}^{\prime};\mathcal{D}_{i}^{\text{query}})$triangleright Outer optimization step
@@ 3.3 Entire Training Pipeline
We follow 3DGS [9] to compute the deviation of the predicted image from the ground truth, using an L1 regularization and a D-SSIM term as the RGB loss. We also employ a sparse loss [8] to encourage the opacity values $\alpha$ of the Gaussian spheres to approach either 00 or $1$, thereby facilitating learning with opaque objects. For normal estimation, we follow GaussianShader [8] to compute the normal vectors using Gaussian’s shortest axis direction, and progressively align them with the depth-inferred pseudo-normals.
@@
We implement MetaGS using PyTorch [17]. For optimization, we use the Adam optimizer [10] with the same parameters as those specified by [9]. A consistent set of hyperparameters is applied across all scenes, with details provided in the Appendix. Our training procedure contains three stages: the first two stages each last for $4k$ iterations, followed by the final meta-learning stage of $5k$ iterations. The entire training process requires roughly one hour on a single NVIDIA RTX 3090 GPU.
@@ 4.1 Experimental Settings
  • Camera-light-colocated relighting: This setup mirrors the one used in IRON [34], where the camera and point light are co-located for each image. Similar to the OOD setting, the colocated configuration also represents a constrained lighting regime, as the light source is fixed to the camera pose. For fair comparison, we generate colocated versions of the synthetic scenes accordingly.
@@
NRHints [33] 17.25 23.92 27.44 18.04 24.63 24.65 22.57 21.55 24.00 23.03
WildLight [5] 21.73 20.95 24.02 18.65 22.53 24.03 21.47 20.33 19.09 20.22
GS3 [1] 18.84 20.30 24.37 17.66 23.34 25.04 21.12 17.34 24.11 19.63
@@
We evaluate MetaGS on $3$ synthetic scenes and $7$ real captured scenes from NRHints [33], featuring a diverse range of materials, complex object shapes, significant self-occlusion, and intricate shadow effects. Each image is rendered (or captured) with a unique camera pose and point-light position. Full details of the datasets are provided in the Appendix.
@@
We compare MetaGS against both NeRF-based and Gaussian-based models, with a particular focus on the OOD relighting setup. The baseline models include state-of-the-art approaches from the past two years: GS3 [1], NRHints [33], WildLight [5], Relightable 3DGS  [7], GaussianShader  [8], and IRON [34]. The evaluation metrics include PSNR, SSIM [24], and LPIPS [37].
@@ 4.3 Generalization to Other Relighting Setups with Constrained OLAT Training
We evaluate MetaGS in the camera-light-colocated setup, as introduced by IRON [34]. Like the OOD scenario, the colocated setup also imposes limited lighting diversity, as the point light is tied to the camera’s position throughout training. Quantitative results in Table 4 show that MetaGS outperforms IRON, achieving more accurate illumination inference under new point positions, and generating fewer rendering artifacts. Visualizations corresponding to these results can be found in the Appendix.
@@
Our model generalizes effectively to environment map relighting. We approximate global illumination by importance sampling point lights from the environment map based on their pixel intensities. Notably, for our model in particular (unlike the compared methods), the sampled light directions at test time may follow a distribution different from those encountered during training, presenting a significant challenge for OOD generalization. Instead of assuming infinite light source distance, we simulate distant lighting by placing point lights along environment directions at a fixed distance (twice the scene radius). We compare our method with R3G [7] and GaussianShader [8]. Both methods fail to converge when trained directly on out-of-distribution (OOD) data, so we construct training data that covers the full hemispherical illumination space. Even with this setup, GaussianShader still fails to converge, while Relightable 3DGS struggles with the OLAT learning task and exhibits noticeable artifacts. As shown in Figure 6, MetaGS produces diverse and visually plausible relighting results under complex environment maps, demonstrating its robustness and generalization capability.
@@
To validate MetaGS’s capability in handling complex lighting conditions, we conduct experiments with test sets containing 22223333 light sources. We also present the results for free-viewpoint relighting with in-distribution point light positions. Please refer to the Appendix for details.
@@ 5 Related Work
Recent differentiable volume rendering techniques, including NeRF-based [16], SDF-based [31, 32], and 3DGS-based [9] methods, have significantly improved the quality and efficiency of novel view synthesis for 3D scenes. NeRF-based methods utilize deep neural networks to model volumetric scene functions, encoding color and density to synthesize high-quality images from sparse viewpoints [16, 14, 36]. However, these approaches typically require substantial training time. 3DGS significantly reduces both training and inference time. It converts point cloud data into a continuous volumetric representation by applying Gaussian kernels to point cloud data, facilitating rendering and further processing. This technique is now widely adopted for efficient 3D and 4D reconstructions across varied data types [27, 8, 15, 25, 30].
The task of 3D relighting involves altering the illumination in a 3D scene while maintaining its geometry. It requires decomposing materials and lighting of a scene from multiple images, which is challenging due to its high-dimensional nature. Recent advances in volume rendering have introduced various solutions for this task, including those based on neural fields [38, 26, 21, 19, 20, 12, 22, 21, 29] and the methods based on Gaussian point clouds [27, 8, 7, 35]. A limitation of these methods is that they require a substantial volume of multi-view images captured under individual lighting conditions, making them impractical in real-world scenarios. Several studies have proposed NeRF-based 3D relighting techniques within the OLAT framework [28, 38, 33], which greatly reduce the demands on training data. Nonetheless, these techniques tend to be computationally intensive because of NeRF’s inherent complexity in the volume rendering process. In parallel, concurrent 3DGS-based approaches [6, 1] have also explored the OLAT relighting challenge. Despite their efforts, these approaches still face difficulties when dealing with lighting conditions not encountered during training, similar to the limitations observed in NeRF-based methods.
@@ Appendix A Notations
      Phong model       $I$       learnable light intensity (global)
@@ Appendix B Datasets
We test on three synthetic scenes and seven real-world scenes from NRHints [33], as shown in Figure 7. Each image is rendered (or captured) with a unique camera pose and point-light position. For each synthetic scene, we generate a total of $600$ OLAT images using Blender, with $500$ for training and $100$ for testing, following either out-of-distribution or colocated data patterns to align with their settings. Specifically, the camera and light source are independently sampled on the upper hemisphere with centered around the scene, with elevation angles ranging from $10^{\circ}$ to $80^{\circ}$, and radial distances uniformly sampled between $4.0$× and $5.0$× the object bounding size. This setup ensures a wide diversity of incident angles while maintaining plausible visibility. In the OOD configuration, training and testing light directions are drawn from non-overlapping hemispheres (azimuth $\phi\in[0,\pi]$ for training; $\phi\in[\pi,2\pi]$ for testing) to induce generalization difficulty. In the colocated setting, camera and light positions are identical, emulating a mobile-light capture scenario. All images are rendered at a resolution of $512\times 512$ using Cycles path tracing in Blender, with $256$ samples per pixel. For real captured scenes in the out-of-distribution experiment, we manually divide the data into training and testing sets, each with lights positioned on opposite sides of the upper hemisphere.
@@
NRHints [33] 17.25 0.8755 0.0872 23.92 0.9485 0.0470 27.44 0.9465 0.0691
WildLight [5] 21.73 0.9217 0.0568 20.95 0.9249 0.0629 24.02 0.9221 0.0789
GS3 [1] 18.84 0.9063 0.0589 20.30 0.9141 0.0628 24.37 0.9349 0.0715
@@
NRHints [33] 18.04 0.7821 0.2405 24.63 0.9289 0.1094 24.65 0.9377 0.0965 22.57 0.8658 0.1396
WildLight [5] 18.65 0.7293 0.2685 22.53 0.8820 0.1448 24.03 0.9246 0.0969 21.47 0.8419 0.1458
GS3 [1] 17.66 0.6919 0.2431 23.34 0.9182 0.1097 25.04 0.9408 0.0954 21.12 0.8264 0.1401
@@
NRHints [33] 21.55 0.8568 0.1435 24.00 0.9279 0.0958 23.03 0.8970 0.1203 22.64 0.8853 0.1351
WildLight [5] 20.33 0.8304 0.1635 19.09 0.8941 0.1129 20.22 0.8330 0.1621 20.90 0.8479 0.1564
GS3 [1] 17.34 0.7684 0.1669 24.11 0.9203 0.0927 19.63 0.8391 0.1329 21.18 0.8436 0.1387

### v3

@@ 1 Introduction
3D scene relighting generates novel lighting effects that interact with the observed 3D environment, with recent advances in learning-based volume rendering offering effective solutions Srinivasan et al. (2021); Zhang et al. (2021a); Toschi et al. (2023); Zheng et al. (2023); Wang et al. (2023); Sun et al. (2023); Chang et al. (2024); Yang et al. (2023). A typical approach captures a substantial number of multi-view images of a scene under individual lighting conditions, then trains a NeRF or 3D Gaussian splatting (3DGS) model to generalize to new point light positions Wu et al. (2024b); Liang et al. (2023); Jiang et al. (2024); Zhu et al. (2024); Gao et al. (2023). Notably, such idealized data requirements are often impractical in real-world scenarios. To simulate a more realistic capturing process, we build on prior work using a One Light At a Time (OLAT) setup Zeng et al. (2023); Liang et al. (2024); Bi et al. (2024), where training data is collected with a moving point light and a moving camera. Recent studies have further explored low-cost OLAT capture using a smartphone flashlight as the moving light source Cheng et al. (2023).
@@
Figure 1: Preliminary results: NRHints Zeng et al. (2023) and GS3 Bi et al. (2024) struggle with out-of-distribution light positions (as specified in Section. 2), as the blurry novel view synthesis results and erroneous normal predictions indicate. GS3 Bi et al. (2024) does not require well-defined surface normals and does not directly output the normal image.
@@
Furthermore, MetaGS integrates simple yet fundamental physical priors into Gaussian splatting by incorporating a learnable Blinn-Phong reflection model Blinn (1977). This approach effectively decouples different shading components—diffuse, specular, and ambient—leading to a more physically grounded understanding of the interactions between objects and lighting in the training data. This design draws inspiration from prior research Zhang et al. (2021b); Wu et al. (2024b); Liang et al. (2023) indicating that modular and compositional representations enhance generalization capabilities. The decoupled rendering improves the model’s ability to generalize to novel lighting conditions, allowing it to independently adjust each shading component according to the specific lighting scenario.
@@ 2 Preliminaries
We summarize existing OLAT relighting methods in Table LABEL:tab:olat_methods. In the preliminary experiments, we evaluate NRHints Zeng et al. (2023), the prior art in NeRF-based relighting models. We observe that it underperforms in testing scenarios with out-of-distribution (OOD) lighting positions, a feature that we believe should be crucial for relighting methods. Specifically, in Figure 1, the lighting in the training set is arranged on one side of the hemisphere, while the lighting in the test set is arranged on the opposite side (cameras are located on both sides). In such cases, NRHints fails to generate reasonable rendering results, likely due to its implicit modeling of shadows and specular reflections. We additionally evaluate a concurrent 3DGS-based approach Bi et al. (2024) and observe similar degraded results under the OOD relighting setup. These findings highlight that understanding intrinsic illumination properties under arbitrary lighting variations remains a challenging task, as the model tends to overfit to perspective-constrained observations and fail to leverage enough physical principles when dealing with unseen lighting distribution.
@@
Zhang et al. (2022) Cheng et al. (2023) Fan et al. (2024) Zeng et al. (2023) Bi et al. (2024) Kuang et al. (2024) Ours
@@ 3.1 Differentiable Phong Model
Our MetaGS method leverages simple yet generalizable physical priors from the Blinn-Phong model Phong (1998), which captures three fundamental components of light transport: ambient, diffuse, and specular reflections. The core idea is to disentangle these illumination components by learning the interactions between (i) the normal vectors of the Gaussian points, (ii) the viewing directions, and (iii) the ray direction from the point light.
@@
where $k_{\{d,s\}}$ and $I_{\{d,s\}}$ are the colors and intensities of the diffuse and specular components. $\lvert\text{lights}\lvert=1$ in OLAT. In graphics, zero-order spherical harmonics (SH) coefficients typically represent the basic, uniform component of a function defined over the sphere, essentially the average or constant part of a lighting environment. Therefore, we restrict Gaussians’ SH coefficients to zero order, denoted by $f_{0}$, corresponding to the ambient color. We compute the diffuse and specular light transports by multiplying the color and intensity of each component, where the intensities $I_{d}$ and $I_{s}$ are defined as:
@@
We calculate shadow effects using a BVH-based ray tracing method Gao et al. (2023). Similar to the camera-to-point accumulated transmittance, the received light intensity $T_{i}^{\text{light}}$ represents the total transmittance of light along the light-to-point ray. It is therefore affected by the opacity of the surfaces encountered along this path. For each point, we determine light visibility by tracing a ray from the Gaussian’s center to the light source. Since only the diffuse and specular colors are influenced by incident light intensity, we update the color of each point by incorporating the light visibility factor into the diffuse and specular terms from Eq. (1):
@@ 3.2 Meta-Learned Gaussian Relighting
To address this, we introduce a meta-learning framework based on bilevel optimization, which has been shown to effectively bridge the distribution shift between the training and testing domains, facilitating the generalization of optimized variables to unseen scenarios Chen et al. (2020). This training strategy can also promote coherent geometry learning in complex OLAT tasks, as shown in the ablation study. In MetaGS, the intuition of incorporating bilevel gradient update is to mitigate overfitting to specific light conditions by explicitly simulating test samples with OOD light sources during each gradient update, thereby improving the model’s ability to generalize to varied lighting scenarios.
@@
1:Input: Training set $\{\mathcal{D}_{t}\}_{1:T}$, where $D_{t}=\{$light position, camera parameters, RGB image}\}
@@
9:  $\{\theta\}\leftarrow\{\theta\}-\beta\sum_{i=1}^{m}\nabla_{\theta}\mathcal{L}_{\text{final}}(\theta_{i}^{\prime};\mathcal{D}_{i}^{\text{query}})$ $\triangleright$ Outer optimization step
@@ 3.3 Entire Training Pipeline
We follow 3DGS Kerbl et al. (2023) to compute the deviation of the predicted image from the ground truth, using an L1 regularization and a D-SSIM term as the RGB loss. We also employ a sparse loss Jiang et al. (2024) to encourage the opacity values $\alpha$ of the Gaussian spheres to approach either 0 or $1$, thereby facilitating learning with opaque objects. For normal estimation, we follow GaussianShader Jiang et al. (2024) to compute the normal vectors using Gaussian’s shortest axis direction, and progressively align them with the depth-inferred pseudo-normals.
@@
We implement MetaGS using PyTorch Paszke et al. (2019). For optimization, we use the Adam optimizer Kingma and Ba (2014) with the same parameters as those specified by Kerbl et al. (2023). A consistent set of hyperparameters is applied across all scenes, with details provided in the Appendix. Our training procedure contains three stages: the first two stages each last for $4k$ iterations, followed by the final meta-learning stage of $5k$ iterations. The entire training process requires roughly one hour on a single NVIDIA RTX 3090 GPU.
@@ 4.1 Experimental Settings
  • Camera-light-colocated relighting: This setup mirrors the one used in IRON Zhang et al. (2022), where the camera and point light are co-located for each image. Similar to the OOD setting, the colocated configuration also represents a constrained lighting regime, as the light source is fixed to the camera pose. For fair comparison, we generate colocated versions of the synthetic scenes accordingly.
@@
NRHints Zeng et al. (2023) 17.25 23.92 27.44 18.04 24.63 24.65 22.57 21.55 24.00 23.03
WildLight Cheng et al. (2023) 21.73 20.95 24.02 18.65 22.53 24.03 21.47 20.33 19.09 20.22
GS3 Bi et al. (2024) 18.84 20.30 24.37 17.66 23.34 25.04 21.12 17.34 24.11 19.63
@@
We evaluate MetaGS on $3$ synthetic scenes and $7$ real captured scenes from NRHints Zeng et al. (2023), featuring a diverse range of materials, complex object shapes, significant self-occlusion, and intricate shadow effects. Each image is rendered (or captured) with a unique camera pose and point-light position. Full details of the datasets are provided in the Appendix.
@@
We compare MetaGS against both NeRF-based and Gaussian-based models, with a particular focus on the OOD relighting setup. The baseline models include state-of-the-art approaches from the past two years: GS3 Bi et al. (2024), NRHints Zeng et al. (2023), WildLight Cheng et al. (2023), Relightable 3DGS  Gao et al. (2023), GaussianShader  Jiang et al. (2024), and IRON Zhang et al. (2022). The evaluation metrics include PSNR, SSIM Wang et al. (2004), and LPIPS Zhang et al. (2018).
@@ 4.3 Generalization to Other Relighting Setups with Constrained OLAT Training
We evaluate MetaGS in the camera-light-colocated setup, as introduced by IRON Zhang et al. (2022). Like the OOD scenario, the colocated setup also imposes limited lighting diversity, as the point light is tied to the camera’s position throughout training. Quantitative results in Table 4 show that MetaGS outperforms IRON, achieving more accurate illumination inference under new point positions, and generating fewer rendering artifacts. Visualizations corresponding to these results can be found in the Appendix.
@@
Our model generalizes effectively to environment map relighting. We approximate global illumination by importance sampling point lights from the environment map based on their pixel intensities. Notably, for our model in particular (unlike the compared methods), the sampled light directions at test time may follow a distribution different from those encountered during training, presenting a significant challenge for OOD generalization. Instead of assuming infinite light source distance, we simulate distant lighting by placing point lights along environment directions at a fixed distance (twice the scene radius). We compare our method with R3G Gao et al. (2023) and GaussianShader Jiang et al. (2024). Both methods fail to converge when trained directly on out-of-distribution (OOD) data, so we construct training data that covers the full hemispherical illumination space. Even with this setup, GaussianShader still fails to converge, while Relightable 3DGS struggles with the OLAT learning task and exhibits noticeable artifacts. As shown in Figure 6, MetaGS produces diverse and visually plausible relighting results under complex environment maps, demonstrating its robustness and generalization capability.
@@
To validate MetaGS’s capability in handling complex lighting conditions, we conduct experiments with test sets containing 2233 light sources. We also present the results for free-viewpoint relighting with in-distribution point light positions. Please refer to the Appendix for details.
@@ 5 Related Work
Recent differentiable volume rendering techniques, including NeRF-based Mildenhall et al. (2021), SDF-based Yariv et al. (2021); Yu et al. (2022), and 3DGS-based Kerbl et al. (2023) methods, have significantly improved the quality and efficiency of novel view synthesis for 3D scenes. NeRF-based methods utilize deep neural networks to model volumetric scene functions, encoding color and density to synthesize high-quality images from sparse viewpoints Mildenhall et al. (2021); Liu et al. (2020); Zhang et al. (2020). However, these approaches typically require substantial training time. 3DGS significantly reduces both training and inference time. It converts point cloud data into a continuous volumetric representation by applying Gaussian kernels to point cloud data, facilitating rendering and further processing. This technique is now widely adopted for efficient 3D and 4D reconstructions across varied data types Wu et al. (2024b); Jiang et al. (2024); Luiten et al. (2024); Wu et al. (2024a); Yang et al. (2024).
The task of 3D relighting involves altering the illumination in a 3D scene while maintaining its geometry. It requires decomposing materials and lighting of a scene from multiple images, which is challenging due to its high-dimensional nature. Recent advances in volume rendering have introduced various solutions for this task, including those based on neural fields Zhang et al. (2021b); Wu et al. (2023); Sun et al. (2023); Sarkar et al. (2023); Srinivasan et al. (2021); Liang et al. (2023); Toschi et al. (2023); Sun et al. (2023); Yang et al. (2023) and the methods based on Gaussian point clouds Wu et al. (2024b); Jiang et al. (2024); Gao et al. (2023); Zhang et al. (2021a). A limitation of these methods is that they require a substantial volume of multi-view images captured under individual lighting conditions, making them impractical in real-world scenarios. Several studies have proposed NeRF-based 3D relighting techniques within the OLAT framework Xu et al. (2023); Zhang et al. (2021b); Zeng et al. (2023), which greatly reduce the demands on training data. Nonetheless, these techniques tend to be computationally intensive because of NeRF’s inherent complexity in the volume rendering process. In parallel, concurrent 3DGS-based approaches Fan et al. (2024); Bi et al. (2024) have also explored the OLAT relighting challenge. Despite their efforts, these approaches still face difficulties when dealing with lighting conditions not encountered during training, similar to the limitations observed in NeRF-based methods.
@@ Appendix A Notations
  Phong model   II   learnable light intensity (global)
@@ Appendix B Datasets
We test on three synthetic scenes and seven real-world scenes from NRHints Zeng et al. (2023), as shown in Figure 7. Each image is rendered (or captured) with a unique camera pose and point-light position. For each synthetic scene, we generate a total of $600$ OLAT images using Blender, with $500$ for training and $100$ for testing, following either out-of-distribution or colocated data patterns to align with their settings. Specifically, the camera and light source are independently sampled on the upper hemisphere with centered around the scene, with elevation angles ranging from $10^{\circ}$ to $80^{\circ}$, and radial distances uniformly sampled between 4.04.0× and 5.05.0× the object bounding size. This setup ensures a wide diversity of incident angles while maintaining plausible visibility. In the OOD configuration, training and testing light directions are drawn from non-overlapping hemispheres (azimuth $\phi\in[0,\pi]$ for training; $\phi\in[\pi,2\pi]$ for testing) to induce generalization difficulty. In the colocated setting, camera and light positions are identical, emulating a mobile-light capture scenario. All images are rendered at a resolution of $512\times 512$ using Cycles path tracing in Blender, with $256$ samples per pixel. For real captured scenes in the out-of-distribution experiment, we manually divide the data into training and testing sets, each with lights positioned on opposite sides of the upper hemisphere.
@@
NRHints Zeng et al. (2023) 17.25 0.8755 0.0872 23.92 0.9485 0.0470 27.44 0.9465 0.0691
WildLight Cheng et al. (2023) 21.73 0.9217 0.0568 20.95 0.9249 0.0629 24.02 0.9221 0.0789
GS3 Bi et al. (2024) 18.84 0.9063 0.0589 20.30 0.9141 0.0628 24.37 0.9349 0.0715
@@
NRHints Zeng et al. (2023) 18.04 0.7821 0.2405 24.63 0.9289 0.1094 24.65 0.9377 0.0965 22.57 0.8658 0.1396
WildLight Cheng et al. (2023) 18.65 0.7293 0.2685 22.53 0.8820 0.1448 24.03 0.9246 0.0969 21.47 0.8419 0.1458
GS3 Bi et al. (2024) 17.66 0.6919 0.2431 23.34 0.9182 0.1097 25.04 0.9408 0.0954 21.12 0.8264 0.1401
@@
NRHints Zeng et al. (2023) 21.55 0.8568 0.1435 24.00 0.9279 0.0958 23.03 0.8970 0.1203 22.64 0.8853 0.1351
WildLight Cheng et al. (2023) 20.33 0.8304 0.1635 19.09 0.8941 0.1129 20.22 0.8330 0.1621 20.90 0.8479 0.1564
GS3 Bi et al. (2024) 17.34 0.7684 0.1669 24.11 0.9203 0.0927 19.63 0.8391 0.1329 21.18 0.8436 0.1387